# Verifiable type-III seesaw and dark matter in a gauged $\boldsymbol{U(1)_{\rm B-L}}$ symmetric model

### v2

@@ V.3 Collider Search
In addition to constraints from relic density and direct or indirect DM searches, the BLB-L gauge sector is subject to stringent experimental limits from collider experiments such as ATLAS, CMS, and LEP-II. Data from LEP-II [57, 58] impose a lower bound on the ratio of the new gauge boson mass to its coupling, requiring it to be at least 7 TeV. However, current LHC experiments have already surpassed the bounds set by LEP-II. In particular, searches for high-mass di-lepton resonances at ATLAS [59] and CMS [60] have placed even tighter constraints on the additional gauge boson, significantly limiting the viable parameter space for the BLB-L extension.
Figure 7: The points satisfying relic density, direct detection, and indirect detection constraints are shown in the $g_{\rm BL}-M_{Z_{\rm BL}}$ plane for the case-I (case-II) with blue circles (green boxes). The constraints from ATLAS[59] and CMS [60] are shown with black and red solid lines, respectively.
To incorporate these collider constraints into our analysis, we follow the approach outlined in [61, 29], where the maximum allowed value of the gauge coupling $g_{\rm BL}$ for a given gauge boson mass $M_{Z_{\rm BL}}$ is determined as
@@
In this section, we briefly discuss the detection prospects of this scenario at collider experiments. Fig. 8 illustrates the production channels of the fermion triplet in proton-proton collisions. We compute the production cross-section of the triplet fermion in proton-proton collisions at a center-of-mass energy of $\sqrt{s}=13$ TeV, shown as a green dotted line in Fig. 9 as a function of $M_{\Sigma}$. As expected, the production cross-section decreases with increasing $M_{\Sigma}$. Existing collider searches at ATLAS [62] and CMS [63] have already placed constraints on the triplet fermion mass, requiring it to be greater than 790 GeV and 880 GeV, as shown in Fig. 9.
Figure 9: Pair production of $\Sigma^{\pm,0}$ at the 13 TeV LHC. The black dotted (solid) line represents the 95% CL upper limit from the CMS [63] (ATLAS [62]) experiment. The green dotted line shows the production cross-section in the absence of $Z_{\rm BL}$. The blue (red) line represents the production cross-section for $M_{Z_{\rm BL}}=5$ TeV and $g_{\rm BL}=0.2$, corresponding to the first (third) generation of triplet fermions.
In the conventional type-III seesaw scenario, the production of $\Sigma$ occurs primarily via SM gauge boson-mediated Drell-Yan processes. However, in a gauged $\rm B-L$ framework, the additional gauge boson $Z_{\rm BL}$ can significantly impact this pair production cross-section. This effect is illustrated in Fig. 9 for benchmark values of $M_{Z_{\rm BL}}=5$ TeV and $g_{\rm BL}=0.2$ consistent with the ATLAS and CMS constraint on additional gauge boson mass and gauge coupling. We consider the production of the first generation of triplets as well as the third generation with the same mass. The third generation of triplets has a B-L charge twice that of the first generation, resulting in a larger cross-section than the latter. Our results show that the presence of $Z_{\rm BL}$ enhances the production cross-section, particularly for $M_{\Sigma}\gtrsim 1.1$ TeV.
@@
Figure 10: A comparison of $\Sigma^{\pm,0}$ pair production cross-sections at the 13 TeV LHC in the minimal type-III seesaw model and its $U(1)_{\rm B-L}$-extended counterpart, focusing on third-generation couplings with $M_{\Sigma}=1.9$ TeV. Regions of parameter space consistent with ATLAS [59] and CMS [60] constraints are indicated by cyan data points. The stars corresponds to three benchmarks as mentioned in Tabel 3.
@@
In Fig. 10, we present the relative enhancement in the production cross-section, defined as ${\sigma(\rm type\text{-}III+U(1)_{\rm B-L})}/{\sigma(\rm type\text{-}III)}$ as a function of $M_{Z_{\rm BL}}$, with the gauge coupling $g_{\rm BL}$ randomly varied within the range $10^{-3},1]$. The figure illustrates that, away from the resonance, the production cross-section can increase by a factor of approximately 50, while near resonance, this enhancement can reach up to 140 times, as indicated by the gradient-colored points.
However, when we impose the ATLAS+CMS constraints on the $g_{\rm BL}-M_{Z_{\rm BL}}$ parameter space, the results are shown by the cyan points. Even after applying these constraints, we still observe an enhancement of up to 40 times, highlighting the significant effect of the $Z_{\rm BL}$ gauge boson in modifying the collider phenomenology of type-III seesaw models. These results highlight the crucial role of the $Z_{\rm BL}$ gauge boson in modifying the collider phenomenology of type-III seesaw models, offering a potential signature for probing this scenario at future high-energy experiments.
@@
Once produced, the charged components $\Sigma_{1}^{\pm}$ predominantly decay via $\Sigma_{1}^{0}\pi^{\pm}$ arising from electroweak radiative corrections that induce a small mass splitting of $\Delta M\simeq 166~\text{MeV}$ between the charged and neutral components of the triplet [32]. This mass splitting enables the two-body pion mode, which accounts for nearly $98\%$ of the total decay width, while the subdominant three-body leptonic channels $\Sigma_{1}^{\pm}\rightarrow\Sigma_{1}^{0}e^{\pm}\bar{\nu}_{e}(\nu_{e})$ and $\Sigma_{1}^{\pm}\rightarrow\Sigma_{1}^{0}\mu^{\pm}\bar{\nu}_{\mu}(\nu_{\mu})$ contribute only $\sim 2\%$. Additionally, due to the mixing of $\Sigma_{1}^{\pm}$ with the SM charged leptons, it can decay through the two body processes such as: $\nu_{l}W^{\pm},l^{\pm}Z,l^{\pm}h,$. The branching ratios of these channels are sensitive to the lightest neutrino mass $m_{1}$ (assuming a normal hierarchy spectrum), as the mixing strength between $\Sigma_{1}$ and the SM leptons depends on the underlying Yukawa structure. On the other hand, the decay channels mediated purely by electroweak corrections, such as $\Sigma_{1}^{\pm}\rightarrow\Sigma_{1}^{0}\pi^{\pm}$ and the leptonic three-body decays mentioned above, remain essentially independent of $m_{1}$ [^2] We note that by setting the rotation matrix $\mathcal{R}=\mathcal{I}_{3\times 3}$ in Eq. (A8), the branching ratio of $\Sigma_{1}^{\pm}$ is found to be highly sensitive to $m_{1}$ in the normal hierarchy (NH) case, while the corresponding ratios for $\Sigma_{2,3}^{\pm}$ remain nearly independent of $m_{1}$ [27]. From here onward, we therefore focus on the collider signatures of $\Sigma_{1}^{\pm}$. A detailed discussion of the decay widths can be found in Appendix A.. In Fig. 11, we present the branching ratios of $\Sigma_{1}^{\pm}$ to different final states for $M_{\Sigma_{1}}=500~\text{GeV}$, assuming a degenerate triplet mass spectrum and $\mathcal{R}=\mathcal{I}$. The best-fit values of neutrino oscillation parameters used in this analysis are listed in Table 2. We observe that for larger $m_{1}$ values, the two-body decays $\nu W^{\pm}$, $\ell^{\pm}Z$, and $\ell^{\pm}h$ dominate, while for smaller $m_{1}\lesssim\mathcal{O}(2\times 10^{-5})~\text{eV}$, the pion channel $\Sigma_{1}^{\pm}\to\Sigma_{1}^{0}\pi^{\pm}$ becomes dominant. In the latter case, $\Sigma_{1}^{\pm}$ behaves as a long-lived charged particle, potentially giving rise to disappearing charged track signatures at colliders.
@@
Figure 12: The total decay length of $\Sigma_{1}^{\pm}$ versus its mass, compared with the ATLAS [64] and CMS [65] bounds on disappearing charged track searches at a center-of-mass energy of 13 TeV.
@@
The ATLAS [64] and CMS [65] searches for disappearing charged tracks have imposed stringent limits on long-lived charged particles such as $\Sigma_{1}^{\pm}$. For instance, current data exclude $\Sigma_{1}^{\pm}$ masses below $640~\text{GeV}$ for $m_{1}=10^{-10}~\text{eV}$, and $606~\text{GeV}$ for $m_{1}=10^{-6}~\text{eV}$. In Fig. 12, we present the decay length of $\Sigma_{1}^{\pm}$ for three choices of $m_{1}:\{5\times 10^{-5},10^{-6},10^{-10}\}$eV as a function of $\Sigma_{1}^{\pm}$ mass. Given that the electroweak-induced mass splitting $\Delta M\simeq 166~\text{MeV}$ naturally leads to a macroscopic decay length, such disappearing track signatures serve as a promising avenue to test this model at present and future collider experiments.
@@
In Fig. 13, we show the exclusion region from ATLAS and CMS experiments in the plane of $m_{1}$ vs $M_{\Sigma_{1}}$. The blue region represent the upper limit on the mass of lightest SM neutrino from Planck data. Here we have used the best fit values of the neutrino oscillation data.
@@
Table 2: The best fit values and the $3\sigma$ ranges of the neutrino oscillation parameters[66].
Figure 14: Effective neutrino mass as a function of lightest neutrino mass for NH (red band) and IH (green band) considering the best fit values of the neutrino oscillation parameters as mentioned in Table 2. Vertical lines represents ATLAS exclusions (regions left to each lines is excluded) for different $\Sigma_{1}$ masses.
The resulting variation of $m_{\beta\beta}$ as a function of $m_{\rm lightest}$ is shown in Fig. 14. The red (green) band corresponds to the NH (IH) spectrum, obtained by scanning over the allowed range of Majorana phases. The gray-shaded region represents the cosmological upper limit on the sum of neutrino masses from the Planck data [1], while the blue-shaded area denotes the exclusion region from KamLAND-Zen [67]. The black dash-dotted vertical lines represent the ATLAS constraints, with the regions to their left excluded for the corresponding values of the heavy fermion mass $M_{\Sigma_{1}}$ in the NH case. We note that similar constraints also arise for the IH scenario, where the contributions are primarily governed by the triplet state $\Sigma_{3}$, and the allowed parameter space depends sensitively on its mass and mixing pattern. From Fig. 14, we observe that the predicted $m_{\beta\beta}$ values may lie within the sensitivity reach of upcoming $0\nu\beta\beta$ experiments. The combined cosmological, collider, and neutrinoless double beta decay bounds thus provide a complementary set of constraints on the model parameter space.
VI Stochastic Gravitational Waves from first-order phase transition
@@
BPs $M_{Z_{\rm BL}}\rm(GeV)$ $g_{{}_{\rm BL}}$ $\lambda$ $u(\rm TeV)$ $v(\rm TeV)$ $T_{C}(\rm TeV)$ $T_{n}(\rm TeV)$ $\alpha$ $\beta/\mathcal{H}$ $M_{\rm DM}(\rm GeV)$ $\sin_{\theta_{h\phi}}$ $\rm Relic$ $\rm DD$ $\rm Collider$ $\Delta\sigma/\sigma\%$
@@
As detailed in Appendix B, we compute the complete effective potential, incorporating the tree-level potential $V_{\rm tree}$, the one-loop Coleman-Weinberg correction $V_{CW}$ [68], and finite-temperature corrections [69, 70]. The critical temperature $T_{c}$, at which the potential develops two degenerate minima $(0,v_{c})$, is determined by analyzing the temperature evolution of the potential. The ratio $v_{c}/T_{c}$ serves as the order parameter, where a larger value indicates a stronger first-order phase transition (FOPT). The FOPT proceeds via quantum tunneling, with the tunneling rate estimated by calculating the bounce action $S_{3}$ following the approach in [71, 72]. The nucleation temperature $T_{n}$ is then determined by equating the tunneling rate to the Hubble expansion rate of the universe, $\Gamma(T_{n})=\mathcal{H}^{4}(T_{n})$.
We then compute the key parameters necessary to estimate the stochastic gravitational wave (GW) spectrum originating from bubble collisions [73, 74, 75, 76, 77], sound waves in the plasma [78, 79, 80, 81], and plasma turbulence [82, 83, 84, 85, 86, 87]. The two critical parameters for estimating the GW signal are the latent heat released relative to the radiation energy density $(\rho_{\rm rad})$ and the duration of the phase transition. These are expressed in terms of $\alpha(T_{n})$ and ${\beta}/{{\mathcal{H}}(T)}$ which determine the strength of the FOPT and are given by:
@@
The details of these contributions are provided in Appendix C.

### v3

@@ 5.3 Collider Search
In addition to constraints from relic density and direct or indirect DM searches, the BLB-L gauge sector is subject to stringent experimental limits from collider experiments such as ATLAS, CMS, and LEP-II. Data from LEP-II Carena:2004xs ; Cacciapaglia:2006pk impose a lower bound on the ratio of the new gauge boson mass to its coupling, requiring it to be at least 7 TeV. However, current LHC experiments have already surpassed the bounds set by LEP-II. In particular, searches for high-mass di-lepton resonances at ATLAS ATLAS:2019erb and CMS CMS:2021ctt have placed even tighter constraints on the additional gauge boson, significantly limiting the viable parameter space for the BLB-L extension.
Figure 7: The points satisfying relic density, direct detection, and indirect detection constraints are shown in the $g_{\rm BL}-M_{Z_{\rm BL}}$ plane for the case-I (case-II) with blue circles (green boxes). The constraints from ATLASATLAS:2019erb and CMS CMS:2021ctt are shown with black and red solid lines, respectively.
To incorporate these collider constraints into our analysis, we follow the approach outlined in Das:2021esm ; Ghosh:2021khk , where the maximum allowed value of the gauge coupling $g_{\rm BL}$ for a given gauge boson mass $M_{Z_{\rm BL}}$ is determined as
@@
In this section, we briefly discuss the detection prospects of this scenario at collider experiments. Fig. 8 illustrates the production channels of the fermion triplet in proton-proton collisions. We compute the production cross-section of the triplet fermion in proton-proton collisions at a center-of-mass energy of $\sqrt{s}=13$ TeV, shown as a green dotted line in Fig. 9 as a function of $M_{\Sigma}$. As expected, the production cross-section decreases with increasing $M_{\Sigma}$. Existing collider searches at ATLAS ATLAS:2020wop and CMS CMS:2019lwf have already placed constraints on the triplet fermion mass, requiring it to be greater than 790 GeV and 880 GeV, as shown in Fig. 9.
Figure 9: Pair production of $\Sigma^{\pm,0}$ at the 13 TeV LHC. The black dotted (solid) line represents the 95% CL upper limit from the CMS CMS:2019lwf (ATLAS ATLAS:2020wop ) experiment. The green dotted line shows the production cross-section in the absence of $Z_{\rm BL}$. The blue (red) line represents the production cross-section for $M_{Z_{\rm BL}}=5$ TeV and $g_{\rm BL}=0.2$, corresponding to the first (third) generation of triplet fermions.
In the conventional type-III seesaw scenario, the production of $\Sigma$ occurs primarily via SM gauge boson-mediated Drell-Yan processes. However, in a gauged $\rm B-L$ framework, the additional gauge boson $Z_{\rm BL}$ can significantly impact this pair production cross-section. This effect is illustrated in Fig. 9 for benchmark values of $M_{Z_{\rm BL}}=5$ TeV and $g_{\rm BL}=0.2$ consistent with the ATLAS and CMS constraints on additional gauge boson mass and gauge coupling. We consider the production of the first generation of triplets as well as the third generation with the same mass. The third generation of triplets has a B-L charge twice that of the first generation, resulting in a larger cross-section than the latter. Our results show that the presence of $Z_{\rm BL}$ enhances the production cross-section, particularly for $M_{\Sigma}\gtrsim 1.1$ TeV.
@@
Figure 10: A comparison of $\Sigma^{\pm,0}$ pair production cross-sections at the 13 TeV LHC in the minimal type-III seesaw model and its $U(1)_{\rm B-L}$-extended counterpart, focusing on third-generation couplings with $M_{\Sigma}=1.9$ TeV. Regions of parameter space consistent with ATLAS ATLAS:2019erb and CMS CMS:2021ctt constraints are indicated by cyan data points. The stars correspond to three benchmarks as mentioned in Table 3.
@@
In Fig. 10, we present the relative enhancement in the production cross-section, defined as
${\sigma(\rm type\text{-}III+U(1)_{\rm B-L})}/{\sigma(\rm type\text{-}III)}$
as a function of $M_{Z_{\rm BL}}$, with the gauge coupling $g_{\rm BL}$ randomly varied within the range $10^{-3},1]$. The figure illustrates that, away from the resonance, the production cross-section can increase by a factor of approximately 50, while near resonance, this enhancement can reach up to 140 times, as indicated by the gradient-colored points.
However, when we impose the ATLAS+CMS constraints on the $g_{\rm BL}-M_{Z_{\rm BL}}$ parameter space, the results are shown by the cyan points. Even after applying these constraints, we still observe an enhancement of up to 40 times. These results highlight the crucial role of the $Z_{\rm BL}$ gauge boson in modifying the collider phenomenology of type-III seesaw models, offering a potential signature for probing this scenario at future high-energy experiments.
@@
We note that the exclusion limits shown in Fig. 9 arise under the assumption that the heavy triplet fermions decay promptly into SM leptons and gauge/Higgs bosons. Such prompt decays require sufficiently large mixing between the triplet and SM leptons, which ensures a decay length much shorter than the detector scale. By contrast, in the regime where the mixing is extremely small, as considered in the following discussion, the charged component $\Sigma^{\pm}_{1}$ becomes long-lived, and the prompt-decay bounds no longer apply. Instead, one must rely on long-lived particle searches (e.g. disappearing charged tracks or displaced vertices).
@@
When the $\Sigma$-leptons mixing is small, once produced, the charged components $\Sigma_{1}^{\pm}$ predominantly decay via $\Sigma_{1}^{0}\pi^{\pm}$ arising from electroweak radiative corrections that induce a small mass splitting of $\Delta M\simeq 166$ MeV between the charged and neutral components of the triplet Cirelli:2005uq . This mass splitting enables the two-body pion mode, which accounts for nearly $98\%$ of the total decay width, while the subdominant three-body leptonic channels $\Sigma_{1}^{\pm}\rightarrow\Sigma_{1}^{0}e^{\pm}\bar{\nu}_{e}(\nu_{e})$ and $\Sigma_{1}^{\pm}\rightarrow\Sigma_{1}^{0}\mu^{\pm}\bar{\nu}_{\mu}(\nu_{\mu})$ contribute only $\sim 2\%$. Additionally, due to the mixing of $\Sigma_{1}^{\pm}$ with the SM charged leptons, it can decay through the two-body processes such as: $\nu_{l}W^{\pm},l^{\pm}Z,l^{\pm}h,$. In general, the mixing between the triplet fermions $\Sigma$ and the SM charged leptons is controlled by the structure of the neutrino Yukawa matrix, which in turn depends on the choice of the orthogonal matrix $\mathcal{R}$ in the Casas–Ibarra parametrization Casas:2001sr . For the special choice $\mathcal{R}=\mathcal{I}_{3\times 3}$, the branching ratios of the two-body decays $\Sigma_{1}^{\pm}\;\to\;\nu_{\ell}W^{\pm},\;\ell^{\pm}Z,\;\ell^{\pm}h$, are sensitive to the lightest neutrino mass $m_{1}$ (assuming normal hierarchy). On the other hand, the decay channels mediated purely by electroweak corrections, such as $\Sigma_{1}^{\pm}\rightarrow\Sigma_{1}^{0}\pi^{\pm}$ and the leptonic three-body decays mentioned above, remain essentially independent of $m_{1}$ [^2] A detailed discussion of the decay widths can be found in A..
In Fig. 11, we present the branching ratios of $\Sigma_{1}^{\pm}$ to different final states for $M_{\Sigma_{1}}=500~\text{GeV}$, assuming a degenerate triplet mass spectrum and $\mathcal{R}=\mathcal{I}$. The best-fit values of neutrino oscillation parameters used in this analysis are listed in Table 2. We observe that for larger $m_{1}$ values, the two-body decays $\nu W^{\pm}$, $\ell^{\pm}Z$, and $\ell^{\pm}h$ dominate, while for smaller $m_{1}\lesssim\mathcal{O}(2\times 10^{-5})~\text{eV}$, the pion channel $\Sigma_{1}^{\pm}\to\Sigma_{1}^{0}\pi^{\pm}$ becomes dominant. In the latter case, $\Sigma_{1}^{\pm}$ behaves as a long-lived charged particle, potentially giving rise to disappearing charged track signatures at colliders.
@@
Figure 12: The total decay length of $\Sigma_{1}^{\pm}$ versus its mass, compared with the ATLAS ATLAS:2022rme and CMS CMS:2018rea bounds on disappearing charged track searches at a center-of-mass energy of 13 TeV.
@@
The ATLAS ATLAS:2022rme and CMS CMS:2018rea searches for disappearing charged tracks have imposed stringent limits on long-lived charged particles such as $\Sigma_{1}^{\pm}$. For instance, current data exclude $\Sigma_{1}^{\pm}$ masses below $640~\text{GeV}$ for $m_{1}=10^{-10}~\text{eV}$, and $606~\text{GeV}$ for $m_{1}=10^{-6}~\text{eV}$. In Fig. 12, we present the decay length of $\Sigma_{1}^{\pm}$ for three choices of $m_{1}:\{5\times 10^{-5},10^{-6},10^{-10}\}$eV as a function of $\Sigma_{1}^{\pm}$ mass. Given that the electroweak-induced mass splitting $\Delta M\simeq 166~\text{MeV}$ naturally leads to a macroscopic decay length, such disappearing track signatures serve as a promising avenue to test this model at present and future collider experiments.
@@
In Fig. 13, we show the exclusion region from ATLAS and CMS experiments in the plane of $m_{1}$ vs $M_{\Sigma_{1}}$. The blue region represents the upper limit on the mass of the lightest SM neutrino from Planck data. Here, we have used the best-fit values of the neutrino oscillation data.
We emphasize that the flavor structure of the heavy–light mixing in the type-III seesaw depends crucially on the choice of the orthogonal matrix $\mathcal{R}$ in the Casas–Ibarra parametrization. While in the earlier part of our analysis we set $\mathcal{R}=\mathcal{I}_{3\times 3}$ for definiteness, this choice corresponds to a particular alignment of the Yukawa couplings and in general need not hold. Relaxing this assumption changes the mixing strength between the triplet fermions and the SM leptons, which in turn will affect the branching ratios and lifetimes of $\Sigma^{\pm}$. Hence conclusions based on $\mathcal{R}=\mathcal{I}$ should be interpreted as representative, not universal.
Figure 14: The total decay length of $\Sigma_{1}^{\pm}$ versus its mass, compared with the ATLAS ATLAS:2022rme and CMS CMS:2018rea bounds on disappearing charged track searches at a center-of-mass energy of 13 TeV considering a general rotation matrix.
In order to check the robustness of our conclusions, we have also performed a numerical scan over a general orthogonal matrix $\mathcal{R}=\mathcal{O}_{23}\,\mathcal{O}_{13}\,\mathcal{O}_{12}$, with complex rotation angles $\theta_{ij}=x_{ij}+iy_{ij}$, varying over a broad range $x_{ij},y_{ij}\in[10^{-4},1]$. Also $m_{1}$ is varied in the range $10^{-10},1$] eV and $M_{\Sigma_{1}}$ in the range $[100,2000]$ GeV. For each choice of $\mathcal{R}$, we recomputed the decay branching ratio of $\Sigma_{1}^{\pm}\rightarrow\Sigma_{1}^{0}\,\pi^{\pm}$ as well as the total decay length. We find that the result is no longer sensitive to $m_{1}$. The result is summarized in Fig. 14, where the color code indicates the branching fraction into the pion channel, and demonstrates that, across a wide span of $\mathcal{R}$, the qualitative conclusion remains valid that in regions where the mixing is small, the pion channel dominates and $\Sigma^{\pm}$ becomes long-lived, yielding disappearing-track signatures at colliders.
@@
Table 2: The best fit values and the $3\sigma$ ranges of the neutrino oscillation parametersdeSalas:2020pgw .
Figure 15: Effective neutrino mass as a function of lightest neutrino mass for NH (red band) and IH (green band) considering the best fit values of the neutrino oscillation parameters as mentioned in Table 2. Vertical lines represent ATLAS exclusions (regions left to each line are excluded) for different $\Sigma_{1}$ masses.
The resulting variation of $m_{\beta\beta}$ as a function of $m_{\rm lightest}$ is shown in Fig. 15. The red (green) band corresponds to the NH (IH) spectrum, obtained by scanning over the allowed range of Majorana phases. The gray-shaded region represents the cosmological upper limit on the sum of neutrino masses from the Planck data Planck:2018vyg , while the blue-shaded area denotes the exclusion region from KamLAND-Zen KamLAND-Zen:2022tow . The black dash-dotted vertical lines represent the ATLAS constraints, with the regions to their left excluded for the corresponding values of the heavy fermion mass $M_{\Sigma_{1}}$ in the NH case. We note that similar constraints also arise for the IH scenario, where the contributions are primarily governed by the triplet state $\Sigma_{3}$, and the allowed parameter space depends sensitively on its mass and mixing pattern. From Fig. 15, we observe that the predicted $m_{\beta\beta}$ values may lie within the sensitivity reach of upcoming $0\nu\beta\beta$ experiments. The combined cosmological, collider, and neutrinoless double beta decay bounds thus provide a complementary set of constraints on the model parameter space.
6 Stochastic Gravitational Waves from first-order phase transition
@@
BPs $M_{Z_{\rm BL}}\rm(GeV)$ $g_{{}_{\rm BL}}$ $\lambda_{\phi^{\prime}}$ $u(\rm TeV)$ $v(\rm TeV)$ $T_{C}(\rm TeV)$ $T_{n}(\rm TeV)$ $\alpha$ $\beta/\mathcal{H}$ $M_{\rm DM}(\rm GeV)$ $\sin_{\theta_{h\phi}}$ $\rm Relic$ $\rm DD$ $\rm Collider$ $\Delta\sigma/\sigma\%$
@@
As detailed in B, we compute the complete effective potential, incorporating the tree-level potential $V_{\rm tree}$, the one-loop Coleman-Weinberg correction $V_{\rm CW}$ Coleman:1973jx , and finite-temperature corrections Dolan:1973qd ; Quiros:1999jp . The critical temperature $T_{c}$, at which the potential develops two degenerate minima $(0,v_{c})$, is determined by analyzing the temperature evolution of the potential. The ratio $v_{c}/T_{c}$ serves as the order parameter, where a larger value indicates a stronger first-order phase transition (FOPT). The FOPT proceeds via quantum tunneling, with the tunneling rate estimated by calculating the bounce action $S_{3}$ following the approach in Linde:1980tt ; Adams:1993zs . The nucleation temperature $T_{n}$ is then determined by equating the tunneling rate to the Hubble expansion rate of the universe, $\Gamma(T_{n})=\mathcal{H}^{4}(T_{n})$.
We then compute the key parameters necessary to estimate the stochastic gravitational wave (GW) spectrum originating from bubble collisions Turner:1990rc ; Kosowsky:1991ua ; Kosowsky:1992rz ; Kosowsky:1992vn ; Turner:1992tz , sound waves in the plasma Hindmarsh:2013xza ; Giblin:2014qia ; Hindmarsh:2015qta ; Hindmarsh:2017gnf , and plasma turbulence Kamionkowski:1993fg ; Kosowsky:2001xp ; Caprini:2006jb ; Gogoberidze:2007an ; Caprini:2009yp ; Niksa:2018ofa . The two critical parameters for estimating the GW signal are the latent heat released relative to the radiation energy density $(\rho_{\rm rad})$ and the duration of the phase transition. These are expressed in terms of $\alpha(T_{n})$ and ${\beta}/{{\mathcal{H}}(T_{n})}$ which determine the strength of the FOPT and are given by:
@@
The details of these contributions are provided in C.